3.1.61 What is the probability that a registered voter voted in the election? About 3.197,707 voted in the election About 3,593.131 of the registered voters did not vote The probability that a registered voter voted in the election is approximately Round to three decimal places as needed.)
Answer #1

Here we have number of registered voter voted in the election = 3197707 and number of registered voter did vote in the election = 3598181

Hence total number of voters = 3197707 + 3598181 = 6795888

We need to find probability of  registered voter voted in the election .

Hence probability = \frac{ number \: of\: registered \: voter \: voted\: in\: the\: election }{Total\: number \: of \: registered \: voters }

  =\frac{3197707}{6795888}

= 0.471